OpenStudy (anonymous):

a. Suppose a market research company finds that at a price of p = $20, they would sell x = 42 tiles each month. If they lower the price to p = $10, then more people would purchase the tile, and they can expect to sell x = 52 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. Hint: Write an equation using two points in the form (x,p). p = -x +62 for p = 20, x = 42 for p = 10, x = 52 A company’s revenue is the amount of money that comes in from sales, before business costs are subtracted. For a single product, you can find

OpenStudy (anonymous):

is ur question cxomplete??

OpenStudy (anonymous):

A company’s revenue is the amount of money that comes in from sales, before business costs are subtracted. For a single product, you can find the revenue by multiplying the quantity of the product sold, x, by the demand equation, p. b. Substitute the result you found from part a. into the equation R = xp to find the revenue equation. Provide your answer in simplified form. The costs of doing business for a company can be found by adding fixed costs, such as rent, insurance, and wages, and variable costs, which are the costs to purchase the product you are selling. The portion of the company’s fixed costs allotted to this product is $300, and the supplier’s cost for a set of tile is $6 each. Let x represent the number of tile sets. c. If b represents a fixed cost, what value would represent b? d. Find the cost equation for the tile. Write your answer in the form C = mx + b. The profit made from the sale of tiles is found by subtracting the costs from the revenue. e. Find the Profit Equation by substituting your equations for R and C in the equation . Simplify the equation. f. What is the profit made from selling 20 tile sets per month? g. What is the profit made from selling 25 tile sets each month? h. What is the profit made from selling no tile sets each month? Interpret your answer

OpenStudy (anonymous):

i. Use trial and error to find the quantity of tile sets per month that yields the highest profit. j. How much profit would you earn from the number you found in part i? k. What price would you sell the tile sets at to realize this profit? Hint: Use the demand equation from part a. 2. The break even values for a profit model are the values for which you earn $0 in profit. Use the equation you created in question one to solve P = 0, and find your break even values.

OpenStudy (dumbcow):

Given p = -x +62 Revenue is price*quantity or px R = x(-x+62) = -x^2 +62x Fixed cost is given as 300 Cost is fixed cost plus variable cost Cost = 6x +300 Profit is Revenue - Cost Profit = (-x^2 + 62x) - (6x+300) Profit = -x^2 +56x - 300 Profit(20) = 420 Profit(25) = 475 Profit(0) = -300 max profit at x=28 Profit(28) = 484 price = -28 +62 = 34